A conductor of length 100cm and area of cross section 1mm^(2) carries a current of 5A .If the resistivity of the material of the conductor is 3×10^(-8)Ohm- m . then the electric field across the conductor is
1
Expert's answer
2020-10-21T09:52:02-0400

Since the voltage is "U=E\times l" , where "E" is the electric field strength and "l" is the length of the conductor (in this case). Since the conductor resistance is "R=\frac{\rho \times l}{S}" , where "\rho" - resistivity, "l" - length of the conductor, and "s" is the cross-sectional area, according to Ohm's law "U=I\times R" , then "E=\frac{U}{l}=\frac{I\times\frac{\rho \times l}{S}}{l}=\frac{5\times 3\times10^{-8} }{1}=15\times10^{-8}" V.